Find the slope and the y intercept of the line 9x=2y+6
Answer by algebrahouse.com(1659)

Slope-intercept form is y = mx + b
m is the slope
b is the y-intercept

9x = 2y + 6
2y = 9x - 6 {subtracted 6 from each side}
y = (9/2)x - 3 {divided each side by 2}

slope = 9/2
y-intercept = -3

For more help from me, visit: www.algebrahouse.com